John decides to lose weight by eating only cold food. A 200 grams mango


pudding yields about 375 kcal of energy when eaten. If its specific heat is 0.5


kcal / kg. ℃, how much greater is its energy content at 60 ℃ than 30 ℃ ?



1
Expert's answer
2021-11-19T10:50:19-0500

"Q=cm\Delta{T}=0.5\frac{kcal}{kg\cdot\degree{C}}\times0.2kg\times(60\degree{C}-30\degree{C})=3\ kcal"


Answer: greater by 3 kcal